OpenStudy (anonymous):

Solve each system of linear equations using substitution. 6x-3y=12 2x-y=6

OpenStudy (anonymous):

FEELS LIKE THAT IS IMPOSSIBLE

OpenStudy (anonymous):

Because if you divide the first equation by 3 then you can get 2x-y=4 but the second equation is 2x-y=6 so 2x-y has two values so it is impossible

OpenStudy (anonymous):

the next one is 2x-5y=4 8x + 12 y=16

OpenStudy (anonymous):

so you can use second equation and divide it by 4 then you can get 2x+3y=4 then you use this one to minus the first equation and get8y=0 so y=0 then you can get X=2

OpenStudy (anonymous):

i was good until the second part

OpenStudy (anonymous):

what do you mean the second part?

OpenStudy (anonymous):

minus the first equation to get 8y=0

OpenStudy (anonymous):

ok,so when you get the equation 2x+3y=4 right? since the first equation is 2x-5y=4 so you use that one to minus the first equation, (2x+3y)-(2x-5y)=4-4, 2x+3y-2x+5y=0, 8y=0, y=0 did you get it know?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem